OpenStudy (anonymous):

5(x + 4) = 7(x - 2) x =

OpenStudy (anonymous):

Use the distibutive property, then solve algebraically

OpenStudy (imtant):

Multiply each number into the brackets 5x+20=7x-14 So 5x-7x=-20-14 x=?

OpenStudy (imtant):

-2x=-34 x=?

OpenStudy (anonymous):

Ill begin it 5(x+4)=7(x-2) 5(x)+5(4)=7(x)-7(2) 5x+20=7x-14

OpenStudy (anonymous):

-34/-2?

OpenStudy (anonymous):

no that'd be irrational. my apologies.

OpenStudy (anonymous):

5x+20=7x-14 -20 -20 ____________________________ 5x=7x-34 -7x -7x ___-__________ -2x=-34 -34/-2= ?

OpenStudy (anonymous):

reduce

OpenStudy (imtant):

Just bring -2 to the right side so it become -34/-2=x x=17

OpenStudy (anonymous):

i got 17!!!!!!!!!!!!

OpenStudy (anonymous):

GJ!

OpenStudy (imtant):

Yes,that's it. :)

OpenStudy (anonymous):

thank you so so much i love you